OpenStudy (anonymous):

What is the sum of a 30–term arithmetic sequence where the first term is 74 and the last term is –100?

OpenStudy (callisto):

probably you can use the formula sum of terms = (a+l)*n/2 where a = first term l = last term n = number of terms i think you've got all the unknowns what you need to do is to substitute the numbers into the formula and get the answer

OpenStudy (amistre64):

posting multiple question on the same topic without interacting with those who are trying to help is a red flag. it indicates that you might simply be posting your homework/test questions to get answers only and not as a means of studying the material to learn. so please feel free to interact with others in the site.

OpenStudy (lgbasallote):

sn = n(a1 + an)/2 a30 = -100 a1 = 74 s30 = 30(74 -100)/2 s30 = 30(26)/2 s30 = 390

OpenStudy (anonymous):

I didm't read the question properly, my apologies. \[15 \times (74-100) = ? \]